A mixture is first made of 150 cm3 of a 0.200 mol.dm–3 solution of silver nitrate and ycm3 of a 2.50 mol.dm–3 solution of potassium cyanide. Thereafter, 100 cm3 of a 0.120 mol.dm─3 solution of potassium chloride is added to this mixture and the resulting solution is diluted to 2.00 dm3 . Calculate the value of y that will just prevent the precipitation of silver chloride.


1
Expert's answer
2021-10-28T07:48:33-0400

150*0.2/100*0.12=30/12=2.5